the average age of 4 brother is
12 years. if tha age of their mother is also
included the average is increased by 5 years
.tha age of the mother is?

Answer by dkppathak(439)

the average age of 4 brother is
12 years. if tha age of their mother is also
included the average is increased by 5 years
.tha age of the mother is?
average age of 4 brothe is 12 yrs
average =total age/number
12=total age /4
total age of 4 brothers =12x4 =48 yrs (let mother age is x)
total age of brothers and mother will be =48+x
new average will be 12+5=17
avrage =toal age/number
17=48+x/5
17x5=48+x
85-48=x
x=37 yrs
mother age is 37 yrs